OpenStudy (anonymous):

A circuit maintains a constant resistance. If the current in the circuit is doubled, what is the effect on the power dissipated by the circuit?

OpenStudy (theeric):

Hi! I would prefer to help you through it. But I don't know where you're stuck. I'll just give you some relevant equivalent expressions: \(P=VI=I^2R=\dfrac {V^2} R\) Two of those will help you. Just look at the quantities mentioned in the problem. Good luck!